OpenStudy (anonymous):

Find the distance between the points (-4, 6) and (8, -12).Round your solution to 2 decimal points.

OpenStudy (anonymous):

\[\sqrt{(8+4)^2+(6+12)^2}\]is what you need to compute

OpenStudy (anonymous):

\[\sqrt{12^2+18^2}=\sqrt{144+324}=\sqrt{466}\] then a calculator

OpenStudy (anonymous):

Thank You Sir??Ma'am?